OpenStudy (anonymous):

Which of the following would produce a negative sum? –8 + 9 –14 + –6 43 + ( –30 )

OpenStudy (anonymous):

@rav can you help please

OpenStudy (moonlitfate):

The second answer. To produce a negative answer, the negative term would have to be bigger than the positive one.

OpenStudy (moonlitfate):

-8 + 9 = 1 ; since 9 is the larger number and it is positive. 43+ (-30) = 13; 13 is the larger number, and is positive.

OpenStudy (anonymous):

-14 + -6 this one

OpenStudy (moonlitfate):

Basically the answer is going to get the sign of the larger number, if that makes sense. :)

OpenStudy (anonymous):

hey can you answer a few more please

OpenStudy (moonlitfate):

Sure! If it will help you understand, I'm all for it.

OpenStudy (anonymous):

Which of the following would produce a positive sum? –12 + 9 15 + ( –19 ) –8 + 10

OpenStudy (anonymous):

Which of the following would produce a positive sum? –5 + 5 28 + ( –4 ) 1 + ( –3 )

OpenStudy (moonlitfate):

Basically, the same rule would apply. So, you can easily go and test all of choices -12 + 9 = -3 15+(-19) = -4 -8+10 = 2

OpenStudy (anonymous):

which one is it

OpenStudy (moonlitfate):

Does it make some sort of sense? Here, try out that last problem that you posted.

OpenStudy (anonymous):

then

OpenStudy (moonlitfate):

Just try the last problem; just getting the answer doesn't help too much. :/

OpenStudy (moonlitfate):

It's better if you understand things.

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

@MoonlitFate The las one will produce a negative answer bc 2 negatives equal a positive

OpenStudy (dmezzullo):

1 negative = negative

OpenStudy (dmezzullo):

So therefore it would be C

OpenStudy (anonymous):

on which one @dmezzullo

OpenStudy (dmezzullo):

the one up top is C

OpenStudy (dmezzullo):

Which of the following would produce a negative sum? –8 + 9 –14 + –6 43 + ( –30 ) this one would be C.

OpenStudy (anonymous):

Which of the following would produce a positive sum? –12 + 9 15 + ( –19 ) –8 + 10 Which of the following would produce a positive sum? –5 + 5 28 + ( –4 ) 1 + ( –3 )

OpenStudy (dmezzullo):

I am not sure about those 2 bc u have to have 2 negatives to get a positive

sam (.sam.):

@dmezzullo 2 in 1 I guess

sam (.sam.):

wait

OpenStudy (dmezzullo):

Wut u mena ?

OpenStudy (dmezzullo):

mean?*

sam (.sam.):

It says positive sum

OpenStudy (anonymous):

@.sam do you know

sam (.sam.):

–8 + 10=2

OpenStudy (dmezzullo):

And then 5+(-5)

OpenStudy (anonymous):

and the other one

OpenStudy (dmezzullo):

A for the second one

OpenStudy (dmezzullo):

I get wut it was asking now(::

sam (.sam.):

I don't think zero is considered positive, so 28 + ( –4 )

OpenStudy (moonlitfate):

@dmezzullo -- Adding two negatives, will give you a negative. :p

OpenStudy (dmezzullo):

No u add 2 negatives and it becomes positive @MoonlitFate

OpenStudy (moonlitfate):

Subtracting is where the changing the sign occurs.

OpenStudy (dmezzullo):

Oh ok

sam (.sam.):

@MoonlitFate unless |-5|+|-3| :P

OpenStudy (dmezzullo):

^

OpenStudy (moonlitfate):

@.Sam. -- Yes, that's true. :3

OpenStudy (anonymous):

so is it this one 28 + ( –4 )

OpenStudy (dmezzullo):

Yes^